OpenStudy (isaiah.feynman):

Rationalize 1/2root3

OpenStudy (isaiah.feynman):

\[\frac{ 1 }{ 2\sqrt{3} }\]

OpenStudy (anonymous):

I could be wrong but wouldn't multiply 2sqrt3? \[\frac{ 1 }{ 2\sqrt{3} }*(\frac{ 2\sqrt{3} }{ 2\sqrt{3} })\]\[=\frac{ 2\sqrt{3} }{ 12 }=\frac{ \sqrt{3} }{ 6 }\]